OpenStudy (horsegal244):

Diane flipped a coin 600 times and it landed heads up 240 times. What is the relative frequency of the coin landing heads up based on the results of this experiment?

OpenStudy (horsegal244):

@Michele_Laino

OpenStudy (anonymous):

Is this multiple question?

OpenStudy (anonymous):

Maybe try dividing

OpenStudy (horsegal244):

ok

OpenStudy (horsegal244):

It is a multiple choice

hartnn (hartnn):

Relative Frequency : How often something happens divided by all outcomes. Example: Your team has won 9 games from a total of 12 games played: the Frequency of winning is 9. the Relative Frequency of winning is 9/12 = 75%

OpenStudy (anonymous):

Okay what are your answers?

OpenStudy (horsegal244):

4% 6% 40% 60%

hartnn (hartnn):

so relative frequency of getting heads is 240/600 and to convert that to %, multiply the answer by 100

OpenStudy (horsegal244):

i dont know how...

hartnn (hartnn):



OpenStudy (horsegal244):

so \[\frac{ 240 }{ 600 }\] = 2.5 or 0.4

hartnn (hartnn):

yesss 0.4 or 0.4*100% = 40%

OpenStudy (horsegal244):

so that is the answer?

hartnn (hartnn):

yep :)